OpenStudy (anonymous):

3y+x>2 and x+2y<=3 what can be said about the value of y?

OpenStudy (anonymous):

3y + x > 2 3y > 2 - x y > (2-x)/3 x + 2y <= 3 2y <= 3 - x y <= (3-x)/2

OpenStudy (anonymous):

3y+x>2 and x+2y<=3 x>2-3y and x<=3-2y then using both equations 3-2y>2-3y i havent used the>= sign because xis greater than 2-3y therefore even if xis equal to 3-2y it will not be equal to 2-3y 3-2>-y y>-1